OpenStudy (anonymous):

what percent of 8 is 5?

OpenStudy (anonymous):

and explanation?

OpenStudy (anonymous):

(5/8)*100 = 500 / 8 = 125 / 2 = 62.5 %

OpenStudy (anonymous):

= 62.5% @Aadarsh is correct :)